OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Jill can mow the same lawn in 4 hours. How many hours would it take them to mow the lawn together?

OpenStudy (anonymous):

bro wtf stop posting up all ur homework problems do them urself

OpenStudy (anonymous):

12/5 hours

OpenStudy (anonymous):

They do a sixth and a quarter per hour or 5/12 -> 12/5 hours.

OpenStudy (anonymous):

whenever you have a problem like this: Someone does something in some amount of time, and someone else does the same something in some other amount of time, this is the formula you want to use to see how long it will take them doing it together: \[\frac{1}{\frac{1}{t_1}+\frac{1}{t_2}}\] Where t1 and t2 are the times for each person.

OpenStudy (anonymous):

Not sure if I did that right, is it 3 hours?

OpenStudy (anonymous):

lets see, it would be: \[\frac{1}{\frac{1}{6}+\frac{1}{4}} = \frac{1}{\frac{4}{24}+\frac{6}{24}} = \frac{1}{\frac{10}{24}} = \frac{24}{10} = \frac{12}{5}\] We have to get that common denominator so we can add those fractions first.

OpenStudy (anonymous):

2.4 hours?

OpenStudy (anonymous):

yes, thats correct :)

OpenStudy (anonymous):

Thanks again :)